OpenStudy (anonymous):

need help with this:4x-2=3x+4...having trouble on how to figure out what the X is

hartnn (hartnn):

try to get like terms together, like 4x and 3x ....4 and -2 to do that first, subtract 3x from both sides, what u get??

OpenStudy (anonymous):

do you mean 3x-4X and then 3x-3X 3x-4 and -2-3X i dont get it

hartnn (hartnn):

subtract 3x means 4x-3x-2=3x-3x+4 did u get this??

OpenStudy (anonymous):

i still dont get it..i am really slow on this and i missed class today so i dont know how t do this at all..honestly...why did you pick the 3x to be on the both sides and not the 4x? can you break it down for me? pretend i am a person that just got is at their first day of school

hartnn (hartnn):

i could select anything, 4x or 3x, both would get me same answer. i selected 3x to be subtracted because, 4x-3x=x, positive. if i would have taken 4x to shift other side,then it would be 3x-4x=-x , negative. dealing with positive is always simpler. its as simple as this: if a+b=c then a=c-b this i got by subtracting b from both sides a+b-b=c-b so a+0=c-b so a=c-b got that?

OpenStudy (anonymous):

x equals 6?

OpenStudy (anonymous):

so bear withme..if i do 4x-3x-2=3x-3x+4 wouldnt it be that the 3x cancel each other out and then the 4x +4-2 equal 6 and the x is 6?

hartnn (hartnn):

@zdstrong yes x=6,but let @idiot23 figure it out.

OpenStudy (anonymous):

@hartnn did I get it?

hartnn (hartnn):

4x-3x-2=3x-3x+4 after this 4x-3x=x 3x-3x=0 so x-2=4 got this much ?

hartnn (hartnn):

then adding 2 to both sides, x-2+2=4+2 so x=6 ok?

OpenStudy (anonymous):

yes, we took the 1st field 4x and subtracted the first of the other side and came up with x, then the next step was to do the same for the fields wihc is 3x-3X which is 0 but i got it- aswer is 6

hartnn (hartnn):

good to hear that :)